# Measuring the splashback feature: Dependence on halo properties and history

### v1

@@ 4 Conclusions
In this work, we have explored the dark matter haloes’ splashback feature dependence on halo physical properties, such as mass, accretion rate and halo formation history, using the MillenniumTNG hydrodynamic simulations. We compare their mutual influence on the halo splashback features and how the splashback features dynamically evolve with time by setting these physical properties as control variables. We have found that the halo splashback features get more prominent with halo mass and time. Quantities that have a strong correlation with halo mass, such as the peak height and concentration, also demonstrate similar influence on the halo splashback depth and width. Through the lens of halo assembly history, we find that splashback width is largely shaped by factors that influence the inner density profile, such as concentration and formation time. Therefore, we provide the analytical fitting equations of the splashback features in terms of these quantities and redshift, aiming to provide a more accurate measure on the halo splashback features. However, the splashback features have less dependence on the instantaneous properties such as accretion rate and the time of halo relaxation. The depth and width record the halo formation’s history instead of capturing the dynamical features at halo boundaries. This complements information from the splashback radius itself. These findings seem to have little dependence on the baryonic physics by comparing the simulation results between the dark matter-only and the hydrodynamic runs of MillenniumTNG with the same resolution. We summarise our results as follows:
  • In Figure 3, we show the depth and width as a function of mass and redshift. The depth increases with mass while the width decreases. We provide a fitting function for these relations in Equations 15 and 3.1
  • In Figure 5, we show the depth and width as a function of accretion rate and redshift. The trend here is less clear than with accretion rate, so we provide a fitting function only at redshift $z=0$ in Equations 17 and 18 for the depth and Equation 19 for the width. We also fit these quantities as a function of redshift for haloes stacked by accretion rate in Equation 21 and 23.
@@
The splashback feature may also be able to probe aspects of a halo that we did not explore here, such as the nature of dark matter. For example, Banerjee et al. (2020) showed that dark matter self interactions causes the slope of the stacked density profile to become slightly shallower compared to CDM, but the position of $R_{\rm sp}$ does not exhibit a significant shift. It may also be possible to further examine the features we did explore by binning stacks by multiple features like accretion rate and formation time to probe the effects of halo history and environment. We present here just an initial study and hope to expand on the more promising features in the future. Projection effects can also affect these features, although effects on the depth and width of the feature can be mitigated with proper fitting (Sun et al., 2025). Observational quantities, like galaxy number density profiles or intracluster light, should also be examined to determine how well they trace these trends in the dark matter haloes. In addition, many of the features, like the merger history or accretion rate, may be affected by stacking haloes rather than studying individual objects. It may be possible to examine objects more individually by stacking by solid angle within a given halo. These unexplored aspects highlight opportunities for future work.
@@ 5 Acknowledgements
We thank Bhuvnesh Jain, Susmita Adhikari, Robyn Sanderson. and Aritra Kundu for helpful discussions. SO is supported by the National Science Foundation under Grant No. AST-2307787. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. The computations are performed on the Engaging cluster supported by the Massachusetts Institute of Technology, the FARSC cluster supported by the Harvard University and the MPCDF cluster supported by the Max Planck Institute.
@@ Appendix A Varying the density profile fitting function
[…]
We stack halos by mass as we did in Section 3.1 and fit the splashback features as a function of mass and redshift. The fitting functions in Equations 15 and 3.1 also apply to the new halo profile in Diemer (2023). The new fitted parameters are (depth: $\chi_{\nu}^{2}=1.4$; width: $\chi_{\nu}^{2}=1.8$):
@@
The consistency of the splashback feature across different density profile models suggests that it is not sensitive to the specific form of the fitting function, but instead represents an intrinsic quantity governed by the physical properties of dark matter haloes. This robustness provides a foundation for the subsequent analyses in the main text, where splashback features are examined as functions of other halo properties beyond mass and redshift.
@@ Appendix B Comparison between the dark matter-only and hydrodynamic simulations
The splashback depth has strongest dependence on the halo mass. We follow the same procedures as described in Section 3.1, stacking halos by mass and fitting the splashback features as functions of mass and redshift using the dark matter-only simulation. The same functional forms defined in Section 2.4, namely Equations 15 and 3.1, are used. For the dark matter-only simulation, the fitted splashback depth and width are (depth: $\chi_{\nu}^{2}=1.4$; width: $\chi_{\nu}^{2}=2.2$):
@@
Similarly, the splashback width has strongest dependence on the peak height, so we bin the haloes by the peak height as we did in Section 3.3.2. We fit the splashback features with the same functional forms as Equations 29, 31 and 32. Using the dark matter-only simulations, they become (depth: $\chi_{\nu}^{2}=3.8$; width: $\chi_{\nu}^{2}=7.5$):
@@
$A_{\mathcal{D}}$ 1.89±0.021.89$\pm$ 0.021.89 ± 0.02 1.97±0.031.97$\pm$ 0.031.97 ± 0.03
$B_{\mathcal{D}}$ 0.44±0.030.44$\pm$ 0.030.44 ± 0.03 0.41±0.030.41$\pm$ 0.030.41 ± 0.03
$C_{\mathcal{D}}$ 0.41±0.04-0.41$\pm$ 0.04- 0.41 ± 0.04 0.49±0.06-0.49$\pm$ 0.06- 0.49 ± 0.06
$D_{\mathcal{D}}$ 1.1±0.11.1$\pm$ 0.11.1 ± 0.1 1.0±0.11.0$\pm$ 0.11.0 ± 0.1
$A_{\mathcal{W}}$ 2.56±0.052.56$\pm$ 0.052.56 ± 0.05 2.5±0.22.5$\pm$ 0.22.5 ± 0.2
$B_{\mathcal{W}}$ 0.87±0.04-0.87$\pm$ 0.04- 0.87 ± 0.04 0.77±0.07-0.77$\pm$ 0.07- 0.77 ± 0.07

### v2

@@ 4 Conclusions
In this work, we have explored the dark matter haloes’ splashback feature dependence on halo physical properties, such as mass, accretion rate and halo formation history, using the MillenniumTNG hydrodynamic simulations. We compare their mutual influence on the halo splashback features and how the splashback features dynamically evolve with time by setting these physical properties as control variables. We have found that the halo splashback features get more prominent with halo mass and redshift. Quantities that have a strong correlation with halo mass, such as the peak height and concentration, also demonstrate similar influence on the halo splashback depth and width. Through the lens of halo assembly history, we find that splashback width is largely shaped by factors that influence the inner density profile, such as concentration and formation time. Therefore, we provide the analytical fitting equations of the splashback features in terms of these quantities and redshift, aiming to provide a more accurate measure of the halo splashback features. However, the splashback features have less dependence on the instantaneous properties, such as accretion rate and the time of halo relaxation. The depth and width record the halo formation’s history instead of capturing the dynamical features at halo boundaries. This complements information from the splashback radius itself. These findings seem to have little dependence on the baryonic physics by comparing the simulation results between the dark matter-only and the hydrodynamic runs of MillenniumTNG with the same resolution. We summarise our results as follows:
  • In Figure 3, we show the depth and width as a function of mass and redshift. The depth increases with mass while the width decreases. We provide a fitting function for these relations in Equations 15 and 3.1.
  • In Figure 5, we show the depth and width as a function of accretion rate and redshift. The trend at higher redshifts is less clear than with mass, so we provide a fitting function only at redshift $z=0$ in Equations 17 and 18 for the depth and Equation 19 for the width. We also fit these quantities as a function of redshift for haloes stacked by accretion rate in Equations 21 and 23.
@@
The splashback feature may also be able to probe aspects of a halo that we did not explore here, such as the nature of dark matter. For example, Banerjee et al. (2020) showed that dark matter self-interactions cause the slope of the stacked density profile to become slightly shallower compared to CDM, but the position of $R_{\rm sp}$ does not exhibit a significant shift. It may also be possible to further examine the features we did explore by binning stacks by multiple features like accretion rate and formation time to probe the effects of halo history and environment. We present here just an initial study and hope to expand on the more promising features in the future. Projection effects can also affect these features, although effects on the depth and width of the feature can be mitigated with proper fitting (Sun et al., 2025). Observational quantities, like galaxy number density profiles or intracluster light, should also be examined to determine how well they trace these trends in the dark matter haloes. In addition, many of the features, like the merger history or accretion rate, may be affected by stacking haloes rather than studying individual objects. It may be possible to examine objects more individually by stacking by solid angle within a given halo. These unexplored aspects highlight opportunities for future work.
@@ 5 Acknowledgements
We thank Bhuvnesh Jain, Susmita Adhikari, Robyn Sanderson, and Aritra Kundu for helpful discussions. SO is supported by the National Science Foundation under Grant No. AST-2307787. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. The computations are performed on the Engaging cluster supported by the Massachusetts Institute of Technology, the FARSC cluster supported by Harvard University and the MPCDF cluster supported by the Max Planck Institute.
@@ Appendix A Varying the density profile fitting function
We fix the parameter $\xi$ at 0.5 as suggested in Diemer (2023), and $R_{200}$ is $1$ in our normalised profiles. When fitting, we follow the suggested ranges in Diemer (2023): $10<\rho_{s}/\rho_{m}<10^{7}$, $0.01<r_{s}/R_{200}<0.45$, $5<r_{t}/R{200}<3, 0.0$$03<\alpha<0.4$, $0.1<\beta<10$, $1<\delta_{1}<100$, $10<\delta_{\rm max}<2000$ and $0.01<s<4$.
We stack halos by mass as we did in Section 3.1 and fit the splashback features as a function of mass and redshift. The fitting functions in Equations 15 and 3.1 also apply to the new halo profile in Diemer (2023). The new fitted parameters are (depth: $\chi_{\nu}^{2}=1.4$; width: $\chi_{\nu}^{2}=1.8$)
@@
The consistency of the splashback feature across different density profile models suggests that it is not sensitive to the specific form of the fitting function but instead represents an intrinsic quantity governed by the physical properties of dark matter haloes. This robustness provides a foundation for the subsequent analyses in the main text, where splashback features are examined as functions of other halo properties beyond mass and redshift.
@@ Appendix B Comparison between the dark matter-only and hydrodynamic simulations
The splashback depth has strongest dependence on the halo mass. We follow the same procedures as described in Section 3.1, stacking halos by mass and fitting the splashback features as functions of mass and redshift using the dark matter-only simulation. The same functional forms defined in Section 2.4, namely Equations 15 and 3.1, are used. For the dark matter-only simulation, the fitted splashback depth and width are (depth: $\chi_{\nu}^{2}=1.4$; width: $\chi_{\nu}^{2}=2.2$)
@@
Similarly, the splashback width has strongest dependence on the peak height, so we bin the haloes by the peak height as we did in Section 3.3.2. We fit the splashback features with the same functional forms as Equations 29, 31 and 32. Using the dark matter-only simulations, they become (depth: $\chi_{\nu}^{2}=3.8$; width: $\chi_{\nu}^{2}=7.5$)
@@
$A_{\mathcal{D}}$ $1.89\pm 0.02$ $1.97\pm 0.03$
$B_{\mathcal{D}}$ $0.44\pm 0.03$ $0.41\pm 0.03$
$C_{\mathcal{D}}$ $0.41\pm 0.04$ $0.49\pm 0.06$
$D_{\mathcal{D}}$ $1.1\pm 0.1$ $1.0\pm 0.1$
$A_{\mathcal{W}}$ $2.56\pm 0.05$ $2.5\pm 0.2$
$B_{\mathcal{W}}$ $0.87\pm 0.04$ $0.77\pm 0.07$